the question says solve the equation 12x + 5 = 14
Answer by checkley77(12844)

12x+5=14
12x=14-5
12x=9
x=9/12
x=3/4 answer.
Proof:
12*3/4+5=14
36/4+5=14
9+5=14
14=14